if the sum of a number and five is tripled, the result is one less than twice the number. find the number????? how do you solve this and write it down
Answer by ankor@dixie-net.com(22740)

if the sum of a number and five is tripled, the result is one less than twice the number. find the number?
:
x = "a number"
:
Just write an expression for each phrase:
:
"sum of a number & 5 is tripled",
3(x+5)
:
"the result"
=
:
"1 less than twice the number"
2x - 1
:
Put is all together to form an equation and solve it:
3(x+5) = 2x - 1
:
3x + 15 = 2x - 1
:
3x - 2x = -1 - 15
:
x = -16
:
:
Now in the original statement see if that is true
"the sum of a number and five is tripled, the result is one less than twice the number"
3(-16+5) = 2(-16) - 1
3(-11) = -32 - 1
-33 = -33; equality confirms our solution of x = -16
:
:
did this shed some light on this for you?